A mass of 1019Kg is supported by two wires A and B each made of a different material. Before loading the wires are of equal length. After loading they are still of equal length. Find the load carried by each wire.
(WA = 3066N) (WB = 6930N)
Wire
Diameter mm
E Nmm-2
A

1

206,000
B
2

116,000
1
Expert's answer
2020-06-10T18:24:40-0400

Solution.

"m=1019kg;\ng=9.81N\/kg;"

"d_A=1mm;"

"d_B=2mm;"

"E_A=206000N\/mm^2;"

"E_B=116000N\/mm^2;"

"F_g=F_A+F_B;"

"F_g=mg;"

"F_g=1019kg\sdot9.81N/kg=9996N;"

"\sigma=\dfrac{F}{S};" "\sigma=E\epsilon \implies F=ES\epsilon;"

"S=\pi \dfrac{d^2}{4};"

"S_A=3.14\sdot\dfrac{(1mm)^2}{4}=0.785mm^2;"

"S_B=3.14\sdot\dfrac{(2mm)^2}{4}=3.14mm^2;"

"F_g=E_AS_A\epsilon+E_BS_B\epsilon;"

"\epsilon=\dfrac{F_g}{E_AS_A+E_bS_B};"

"\epsilon=\dfrac{9996N}{206000N/mm^2\sdot0.785mm^2+116000N/mm^2\sdot3.14mm^2}=0.019;"

"F_A=206000N/mm^2\sdot0.785mm^2\sdot0.019=3072N;"

"F_B=116000N/mm^2\sdot3.14mm^2\sdot0.019=6924N;"

Answer: "F_A=3072N;"

"F_B=6924N."